OpenStudy (anonymous):

integrate 1/(root(1-2x) -root(1-5x)) I don't seem to be able to identify any forms for this... Please help maybe :3

OpenStudy (anonymous):

\[\int \frac{1}{\sqrt{1-2x}-\sqrt{1-5x}} \text{d}x\]right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

we had this for our exam, i still cant figure out a way to integrate it...

OpenStudy (anonymous):